# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@ Appendix B Current and density profiles in the system approaching steady state
In this Appendix we seek to compute the time evolution of the current produced by a periodic TASEP starting from a clustered configuration, i.e all particles being adjacent one to the other. It has been proven [38] that the evolution of the density profile, seen as a continuous variable in time and space, in the TASEP is given by the unique entropic weak solution of a generalized Burgers’ equation with appropriate boundary and initial conditions. From now onward, to avoid ambiguity, we will use $\overline{\rho}$ to refer to the number of particles divided by the lattice length, while for the density profile in the continuum limit we will use $\rho(x,t)$. Thus, in the case of a clustered initial state and periodic boundaries on has:
@@
Figure 11: Comparison between the simulated and predicted current as a function of time, starting from a configuration with all particles clustered on the first sites. The horizontal line correspond to the expected steady-state value.
@@
Starting from $t=\overline{\rho}L$ the $0-$density characteristics will intersect with the rarefaction fan’s characteristics producing a shock, which will itself exist up until the last 0-characteristic will have intersected it. Depending on the starting cluster length the shock may or may not cross the system boundary, in which case its periodicity has to be accounted for. The shock curve $s(t)$ can be computed by solving the ODE (22) along with the condition $x(\overline{\rho}L)=0$, leading to:
@@
To understand where this equation is defined we have to compute the intersection point $x^{*}$ with $x=t-(1-\rho)L$ (the last $0-$characteristic) in non-periodic space and then check how many lattice lengths are spanned by the two curves before intersecting. We obtain $x^{*}=\left(\frac{1}{4\overline{\rho}}-1+\overline{\rho}\right)L$, as such the shock will cross the boundary $n=\left[\frac{x^{*}}{L}\right]=\left[\frac{1}{4\overline{\rho}}-1+\overline{\rho}\right]$ times, where the square brackets mean that we are considering the integer part of their argument. Thus we obtain:
@@
for $t\in\left[(1-\overline{\rho})L+(i-1)L,(1-\overline{\rho})L+iL\right)$ Finally, we get to the time instant $t^{*}=\frac{L}{4\overline{\rho}}$ where the last $0-$characteristic line meets the shock:
@@
Figure 11 shows how the current $J_{0}$, computed in Eq. (16) after plugging $\rho(x,t)$ obtained in this Appendix into Eq. (14). We notice that the TASEP steady-state value given in Eq. (1) is reached in a short time compared to the duration of the state with zero paused particles given by $k_{+}\rho L$. This further justifies approximating the current in this state with Eq. (1), as done in the multi-cluster approach (Section IV.2.3).

### v2

@@ Appendix B Current and density profiles in the system approaching steady state
In this Appendix we seek to compute the time evolution of the current produced by a periodic TASEP starting from a single cluster configuration, i.e all particles being adjacent one to the other. It has been proven [39] that the evolution of the density profile, seen as a continuous variable in time and space, in the TASEP is given by the unique entropic weak solution of a generalized Burgers’ equation with appropriate boundary and initial conditions. From now onward, to avoid ambiguity, we will use $\overline{\rho}$ to refer to the number of particles divided by the lattice length, while for the density profile in the continuum limit we will use $\rho(x,t)$. Thus, in the case of a single cluster initial state and periodic boundaries on has:
@@
Figure 11: Comparison between the simulated and predicted current as a function of time, starting from a configuration with all particles clustered on the first sites. The horizontal line corresponds to the expected steady-state value.
@@
Starting from $t=\overline{\rho}L$ the 00-0 -density characteristics will intersect with the rarefaction fan’s characteristics producing a shock, which will itself exist up until the last 0-characteristic will have intersected it. Depending on the starting cluster length the shock may or may not cross the system boundary, in which case its periodicity has to be accounted for. The shock curve $s(t)$ can be computed by solving the ODE (22) along with the condition $x(\overline{\rho}L)=0$, leading to:
@@
To understand where this equation is defined we have to compute the intersection point $x^{*}$ with $x=t-(1-\rho)L$ (the last 00-0 -characteristic) in non-periodic space and then check how many lattice lengths are spanned by the two curves before intersecting. We obtain $x^{*}=\left(\frac{1}{4\overline{\rho}}-1+\overline{\rho}\right)L$, as such the shock will cross the boundary $n=\left[\frac{x^{*}}{L}\right]=\left[\frac{1}{4\overline{\rho}}-1+\overline{\rho}\right]$ times, where the square brackets mean that we are considering the integer part of their argument. Thus we obtain:
@@
for $t\in\left[(1-\overline{\rho})L+(i-1)L,(1-\overline{\rho})L+iL\right)$ Finally, we get to the time instant $t^{*}=\frac{L}{4\overline{\rho}}$ where the last 00-0 -characteristic line meets the shock:
@@
Figure 11 shows how the current $J_{0}$, computed in Eq. (16) after plugging $\rho(x,t)$ obtained in this Appendix into Eq. (14). We notice that the TASEP steady-state value given in Eq. (1) is reached in a short time compared to the duration of the state with zero paused particles given by $k_{p}\rho L$. This further justifies approximating the current in this state with Eq. (1), as done in the multi-cluster approach (Section IV.2.3).